# Does Mapo Tofu Contain Coffee? Probing LLMs for Food-related Cultural Knowledge

### v1

@@ Abstract
[…]
@@ 1 Introduction
Trained on vast datasets, Large Language Models (LLMs) encode a wide array of knowledge, but also face challenges with various biases, such as those related to gender Savoldi et al. (2021); Kaneko et al. (2022), belief Søgaard (2021); González et al. (2021); Lent and Søgaard (2021), and culture Cao et al. (2023); Deshpande et al. (2022); Yin et al. (2022). In this paper, we concentrate on probing LLMs for cultural knowledge, aiming to provide evidence of cultural biases inherent in LLMs and explore potential strategies for reducing these cultural biases. Specifically, we begin by introducing the FmLAMA construction method (§3), followed by the knowledge probing method (§4), encompassing the probing template, task, and metric. Subsequently, we conduct experiments employing English prompts on the filtered sub-dataset FmLAMA-en to probe both encoder-only and encoder-decoder LLMs (§5.1). We then analyze various aspects of probing prompt settings (§5.2), including cultural background references, code-switching settings, and language choices within prompts. Following this, we proceed with additional probing experiments targeting decoder-only LLMs (§5.3), considering the prevalence of current advanced LLMs. Finally, we delve into case studies (§5.4) and human evaluations(§6). Figure 1 provides an overview of our work’s various aspects, and our contributions are:
  • We present FmLAMA, a pioneering dataset focused on the food domain, which is inherently rich in cultural diversity Cao et al. (2024). This dataset is a multifaceted tool for probing LLMs across cultures and languages.
  • We propose novel metrics designed to assess LLMs’ ability to accurately and sensitively probe for cultural knowledge, incorporating both absolute- and fuzzy-match techniques.
@@ 2 Related Work
Cultural knowledge, encompassing the customs, beliefs, traditions, and practices of a culture, is crucial yet challenging to encapsulate. While some researchers focus on manually curating cultural knowledge datasets, others evaluate LLMs’ performance on culturally related tasks. Yin et al. (2022) and Palta and Rudinger (2023) have developed benchmarks such as geo-diverse prompts and food-custom datasets (FORK) to probe cultural biases in commonsense reasoning systems. However, manual dataset construction is inefficient and hard to scale, prompting a shift towards automated methods. For instance, StereoKG Deshpande et al. (2022) offers a scalable knowledge graph that blends cultural knowledge with stereotypes, and Candle Nguyen et al. (2023) extracts cultural commonsense knowledge from the web, organizing it into clusters. Despite these advances, the variability in data representation—from sentences to triplets using OpenIE—poses challenges for consistency and noise control in knowledge probing. Keleg and Magdy (2023) aims to mitigate this by selecting culturally diverse factual triples from Wikidata, focusing mainly on explicit country information. In contrast, our work proposes an automated, efficient approach to constructing a cultural knowledge dataset in a uniform triplet format, addressing the limitations of existing methods and focusing on implicit cultural knowledge. Table 1 contrasts our dataset with prior cultural knowledge collections.
Datasets Format Topic Construction method Size
FmLAMA (ours) Triplet knowledge Food domain Automatically constructed 33601
@@
Deciphering the knowledge encoded by LLMs poses significant challenges due to their opaque nature , early benchmarks like LAMA Petroni et al. (2019) sought to quantify the factual knowledge in English LLMs, while ParaRel Elazar et al. (2021) highlighted their consistency issues. Subsequent efforts as mLAMA Kassner et al. (2021) and mParaRel Fierro and Søgaard (2022) expanded these benchmarks multilingually, though such methods often focus on single-word entities, limiting their depth of assessment. To address these shortcomings, newer studies Shin et al. (2020); Zhong et al. (2021); Meng et al. (2022) have evolved towards eliciting more comprehensive factual knowledge, including multi-word entities, with Jiang et al. (2020a) developing algorithms for multi-token predictions. LPAQA Jiang et al. (2020b) further refines this by optimizing prompt discovery for more accurate knowledge probing. Our work builds on this foundation, targeting multi-token probing within the food domain, characterized by complex expressions like Trigonella foenum-graecum and almond paste. We also introduce absolute- and fuzzy-match metrics for a nuanced evaluation of LLMs’ cultural knowledge.
@@ 3 FmLAMA Construction
iii. Properties included. We prioritize the property “has part(s)” to identify food ingredients for each dish. Additional properties like “made from material” and “image” are collected to support future research (e.g., multimodal), though they are not utilized in this study. Language consistency for property descriptions ensures uniformity across the dataset. Our dataset, FmLAMA, comprises 33,601 dish instances, detailed by name, origin, ingredients, and optionally, materials and images. Examples are provided in Appendix A.
Step #3: Filter by language.
To explore various language settings, we further create sub-datasets by applying language filtering to the created FmLAMA dataset. This results in FmLAMA-la, where ‘la’ designates the specific language of the current sub-dataset. In this paper, We focus on a typologically diverse set of languages, namely, English (en), Chinese (zh), Arabic (ar), Korean (ko), Russian (ru), and Hebrew (he), with a filtered sub-dataset of 2590, 815, 571, 807, 961, 462 dishes each. These languages span 4 different language families – Indo-European (English, Russian), Semitic (Hebrew, Arabic), Altaic (Korean), and Sino-Tibetian (Chinese), and are spoken by more than 2.356 billion speakers. Furthermore, these languages represent cultural diversity, being spoken on different continents by groups with rich and distinct cultural backgrounds.
@@ 4 Cultural Knowledge Probing
4.1 Probing templates
Following most knowledge-probing methods, we adopt Word Predictions (WP) as knowledge-probing tasks. Specifically, we manually design prompt templates focused on the core attribute “has part(s)”, illustrating a connection between a dish (subject) and its ingredient(s) (object). Considering LLMs produce varied predictions based on prompt framing Elazar et al. (2021); Wang et al. (2023), we craft five templates conveying identical meanings in each prompt language. These templates, in six various involved languages, are depicted in Figure 7 in Appendix B.
Cultural context
To explore the impact of introducing cultural context on LLMs’ ability to access cultural knowledge, we enhance the basic templates by integrating location adverbials. For instance, “In [C], the food [X] includes the ingredients of [Y]”, where [C] denotes the country of origin for the dish [X], [Y] indicates the ingredient object.
Code-switching
To assess the cultural knowledge probing ability of multilingual LLMs under code-switching settings, we configure prompt variations with English-Chinese and Chinese-English CS settings, specifically based on purely English and Chinese prompts, as follows:
  1. EN_en:
    The food beef bourguignon has the ingredients of [Y].
  2. EN_zh:
    The food 红酒炖牛肉 has the ingredients of [Y].
  3. ZH_zh:
    美食红酒炖牛肉是以[Y]为主要原料制作的。
  4. ZH_en:
    美食beef bourguignon是以[Y]为主要原料制作的。
Here, EN and ZH signify the main language of the prompt, while en and zh indicate the subject’s language.
@@ 4.2 Probing task
The probing task is defined as a candidate object retrieval problem. The candidate set consists of all objects in the filtered sub-dataset FmLAMA-la. [^3] As the size of the filtered sub-dataset increases, the candidate object set also expands, leading to greater difficulty in probing. Consequently, in this paper, results obtained by probing across different filtered sub-datasets cannot be used for horizontal comparison. Our primary objective is to utilize LLMs to obtain the probability of each candidate $C$ and subsequently rank the predicted objects based on these probabilities.
@@
We use the corresponding subject-object tuples ([X], [Y]) as the query and probe LLMs by replacing the subject and masking the object. Considering that each candidate object is tokenized into $d$ subtokens $\left\{c_{1},\cdots,c_{d}\right\}$ by LLMs correspondingly, we apply [MASK] token of varying lengths to the objects within each query. So we construct $D$ queries for each food subject $X$ based on the same template $t$, $D$ is the maximum number of object tokens, each query $Q_{d}^{X}$ is defined as:
$Q_{d}^{X}=t\left(X,\left[\mathrm{MASK}\right]*d\right)\vspace{-3pt}.$ (1)
@@
We use Mean Pooling [^4] Mean Pooling is usually used in multi-token probing and much better than max-pooling and first-pooling methods, this is demonstrated by the experimental results of different pooling methods from  Wu et al. (2023). method to obtain the prediction probability of each candidate. Specifically, for candidate object $C=\left\{c_{1},\cdots,c_{d}\right\}$ of length $d$, we obtain its probability from the likelihoods associated with the [MASK] tokens in $Q_{k}^{X}$, and the predicted probability of $C$ is calculated as the average of the probabilities of composing its subtokens:
$P\left(Q_{d}^{X},C\right)=\frac{1}{d}\sum_{i=1}^{d}{p\left(Q_{d}^{X},\left[\mathrm{MASK}\right]_{i}==c_{i}\right)}\vspace{-3pt},$ (2)
where $p\left(\cdot\right)$ is obtained after the log softmax operation.
@@ 4.3 Probing Metric
mAP is widely used in information retrieval settings, assessing the relevance of predicted objects (in our case ingredients) only when they precisely match the golden ones. The precision at rank $k$ (P@k) for a given food instance $i$ is defined as follows:
@@
where $\mathrm{ing}_{i}$ is the golden ingredients set, $\mathrm{topk}_{i}$ signifies the set of top-$k$ objects with the highest predicted probability of belonging to food item $i$ by LLMs. Then the average precision of food item $i$ is computed as follows:
$\mathrm{AP}_{i}=\frac{1}{\left|\mathrm{ing}_{i}\right|}\sum_{k}^{n}{\mathrm{P}@\mathrm{k}\times\mathrm{rel}@\mathrm{k}}\vspace{-3pt},$ (4)
where $n$ refers to the size of the candidate object set and $\mathrm{rel}@\mathrm{k}$ is a relevance indicator function, which equals 1 if the object at rank $k$ is relevant to food item $i$ and equals to 0 otherwise. Finally, we compute the mAP in the following way:
@@
where G represents a food group we are focusing on (i.e. a subset of FmLAMA).
Origin Count Encoder-only LLMs Encoder-Decoder LLMs Avg.
Bb-c Bb-u Bl-c Bl-u mB-c mB-u mT5 T5
Italy 215 (8.3%) 3.93±2.29 5.43±2.26 4.39±2.44 5.32±3.14 5.25±2.40 5.65±4.03 4.49±1.37 5.09±1.82 4.94
U.S. 285 (11.0%) 10.26±4.94 14.58±6.29 11.60±4.67 12.72±6.25 13.79±5.16 13.10±7.62 12.37±3.37 13.16±5.32 12.70
Turkey 98 (3.8%) 7.80±5.13 10.53±5.09 6.72±3.97 8.02±5.66 11.55±6.37 9.12±6.28 6.29±2.77 7.84±3.98 8.48
Japan 186 (7.2%) 7.99±3.75 7.53±4.09 8.68±4.03 6.64±3.75 7.18±3.41 7.59±4.59 5.57±2.63 5.64±1.40 7.10
France 175 (6.8%) 5.01±2.99 6.57±2.90 6.15±2.58 5.55±2.29 5.64±2.62 6.02±3.59 3.63±1.90 4.05±1.67 5.33
U.K. 83 (3.2%) 8.40±5.34 8.84±3.39 11.17±6.01 9.29±5.83 10.48±4.16 8.52±6.48 6.29±3.94 7.71±3.55 8.83
Mexico 57 (2.2%) 6.41±3.12 6.93±2.95 7.72±2.58 7.16±4.55 8.08±2.64 7.92±2.26 3.61±0.97 4.18±2.72 6.50
India 132 (5.1%) 12.63±6.56 14.18±5.99 13.37±7.32 12.10±6.99 9.78±4.34 10.56±5.67 8.64±3.41 5.41±2.55 10.83
Germany 57 (2.2%) 4.94±3.21 5.42±3.39 5.64±3.43 5.78±2.69 7.40±2.65 7.81±4.92 4.20±1.61 5.79±1.26 5.87
China 97 (3.8%) 10.43±4.17 12.35±3.13 12.36±3.34 11.27±4.69 12.54±3.47 11.79±5.54 7.93±1.97 9.36±3.18 11.00
Iran 21 (0.8%) 7.00±5.17 5.96±4.26 6.32±4.56 6.33±3.82 8.31±4.73 9.42±5.85 10.37±3.09 5.67±0.42 7.42
Greece 21 (0.8%) 3.15±2.93 3.99±2.46 2.78±1.84 2.73±1.63 3.72±3.24 3.12±1.42 2.12±0.92 0.73±0.27 2.79
Spain 95 (3.7%) 3.54±1.93 5.92±2.74 4.26±2.24 5.37±2.89 4.61±1.84 5.42±1.90 3.26±1.77 3.01±1.17 4.42
Russia 27 (1.0%) 5.55±3.53 7.68±2.17 5.34±3.21 5.48±3.64 5.31±3.22 5.57±2.58 1.45±1.28 1.85±0.23 4.78
Others 1031 (40.0%) 5.92±3.04 6.30±2.81 6.41±2.88 6.37±3.17 6.06±2.51 5.62±2.42 3.82±1.21 3.61±1.06 5.51
ALL 2580 (100.0%) 6.87±3.46 8.09±3.40 7.59±3.36 7.51±3.77 7.70±2.91 7.43±3.73 5.51±1.75 5.58±1.95 7.04
(a) Performance results evaluated using mAP (%).
Origin Encoder-only LLMs Encoder-Decoder LLMs Avg.
Bb-c Bb-u Bl-c Bl-u mB-c mB-u mT5 T5
Italy 0.3107±0.07 0.3391±0.05 0.3206±0.07 0.3246±0.05 0.2992±0.05 0.3279±0.04 0.3287±0.02 0.3097±0.01 0.3201
U.S. 0.3593±0.08 0.3957±0.07 0.3793±0.07 0.3813±0.07 0.3730±0.05 0.3760±0.06 0.3796±0.03 0.3844±0.04 0.3786
Turkey 0.3431±0.07 0.3592±0.08 0.3542±0.07 0.3521±0.06 0.3478±0.08 0.3480±0.05 0.3501±0.04 0.3243±0.03 0.3474
Japan 0.3247±0.07 0.3321±0.06 0.3312±0.08 0.3072±0.05 0.2902±0.05 0.2965±0.05 0.3072±0.03 0.2815±0.01 0.3088
France 0.3280±0.07 0.3420±0.04 0.3348±0.05 0.3372±0.03 0.3261±0.05 0.3278±0.03 0.3342±0.02 0.3268±0.02 0.3321
U.K. 0.3132±0.08 0.3343±0.03 0.3401±0.06 0.3362±0.05 0.3177±0.04 0.3255±0.05 0.3126±0.04 0.3079±0.02 0.3234
Mexico 0.3123±0.05 0.3418±0.03 0.3473±0.06 0.3254±0.05 0.3087±0.05 0.3152±0.02 0.3276±0.01 0.3130±0.01 0.3239
India 0.3638±0.09 0.3741±0.08 0.3796±0.09 0.3531±0.08 0.3215±0.06 0.3367±0.07 0.3352±0.05 0.2920±0.02 0.3445
Germany 0.3127±0.10 0.3268±0.08 0.3362±0.08 0.3191±0.06 0.3118±0.05 0.3473±0.06 0.3189±0.05 0.3263±0.03 0.3249
China 0.3210±0.04 0.3493±0.02 0.3476±0.02 0.3354±0.03 0.3398±0.02 0.3309±0.04 0.3569±0.02 0.3156±0.03 0.3371
Iran 0.3195±0.07 0.3145±0.09 0.3399±0.08 0.3145±0.08 0.3252±0.07 0.3476±0.04 0.3555±0.03 0.3233±0.04 0.3300
Greece 0.3113±0.08 0.3399±0.06 0.3359±0.08 0.3178±0.06 0.2822±0.08 0.3180±0.02 0.3530±0.05 0.3048±0.01 0.3204
Spain 0.3008±0.06 0.3280±0.05 0.3189±0.06 0.3117±0.04 0.2856±0.05 0.3068±0.03 0.3168±0.02 0.2828±0.01 0.3064
Russia 0.3368±0.05 0.3680±0.04 0.3355±0.06 0.3397±0.04 0.3113±0.07 0.3158±0.02 0.3313±0.03 0.2721±0.02 0.3263
Others 0.3235±0.06 0.3378±0.04 0.3376±0.06 0.3263±0.04 0.3088±0.05 0.3150±0.02 0.3266±0.02 0.2821±0.02 0.3197
ALL 0.3278±0.06 0.3468±0.05 0.3428±0.06 0.3339±0.05 0.3177±0.05 0.3264±0.03 0.3338±0.02 0.3050±0.02 0.3293
(b) Performance results evaluated using mWS.
Table 2: Probing performance comparison with English prompts and FmLAMA-en sub-dataset. “B/mB” respectively represent abbreviations for BERT and mBERT. “b/l” stands for base/large and “c/u” stands for cased/uncased. Bold represents the best-performing cultural group within the same model (each column). Red indicates the best-performing LLMs in each specific cultural group (each row). We find that LLMs typically score higher on U.S. cultural knowledge, and monolingual English LLMs perform better in English-speaking countries. The Pearson correlation coefficient between mAP and mWS is 0.66 based on the results above.
@@
mWS is defined based on the semantic similarity between predicted and golden objects. First, we define the similarity score $S\left(i,g\right)$ for each ingredient $g$ within each food instance $i$. Only the predicted objects in the top-$l$ rankings that are most similar to $g$ contribute to the evaluation score, where $l$ is the size of $\mathrm{ing}_{i}$. Mathematically, $S\left(i,g\right)$ is defined as follows:
@@
where $w_{g}$ and $w_{p}$ are the continuous word representations for the ingredient $g$ and the predicted object $p$, respectively, and are obtained using Fasttext Bojanowski et al. (2017); Joulin et al. (2016), $\cos\left(\cdot\right)$ is the cosine similarity function. Then we can compute the probing similarity $WS_{i}$ for each food instance $i$ and mWS for the targeted food group as follows:

### v3

@@ Abstract
Li Zhou1,2, Taelin Karidi3, Wanlong Liu4, Nicolas Garneau5, Yong Cao6, Wenyu Chen4, Haizhou Li1,2, Daniel Hershcovich5 1The Chinese University of Hong Kong, Shenzhen 2Shenzhen Research Institute of Big Data 3Hebrew University of Jerusalem 4University of Electronic Science and Technology of China, 5University of Copenhagen, 6University of Tübingen, Tübingen AI Center {lizhou21, haizhouli}@cuhk.edu.cn, taelin.karidi@mail.huji.ac.il, liuwanlong@std.uestc.edu.cn, {nicolas.garneau, dh}@di.ku.dk, yongcao2018@gmail.com, cwy@uestc.edu.cn
@@ 1 Introduction
Trained on vast datasets, Large Language Models (LLMs) encode a wide array of knowledge, but also face challenges with various biases, such as those related to gender Savoldi et al. (2021); Kaneko et al. (2022), belief Søgaard (2021); González et al. (2021); Lent and Søgaard (2021), and culture Cao et al. (2023); Deshpande et al. (2022); Yin et al. (2022); Mukherjee et al. (2024); Singh et al. (2024). In this paper, we adopt food-domain knowledge data as cultural semantic proxy, “country of origin” as cultural demographic proxy Adilazuarda et al. (2024), and address the following research questions: (1) What cultural biases exist within the cultural knowledge contained in LLMs? (2) How can we best prompt and evaluate LLMs to elicit culture-specific knowledge? Specifically, we first design an automated construction method for the dataset FmLAMA3). Next, we introduce the cultural knowledge probing method (§4), encompassing the probing task, template and metric. We then conduct experiments (§5) and analysis (§6) to address our proposed research questions. Finally, we perform an error analysis on the probing results (§7). Figure 1 provides an overview of our work’s various aspects, and our contributions are:
  • We present FmLAMA, a pioneering dataset focused on the food domain, which is inherently rich in cultural diversity. This dataset is a multifaceted tool for probing LLMs across cultures and languages.
  • We propose novel metrics aimed at evaluating LLMs’ capacity to accurately and sensitively probe cultural knowledge, utilizing a combination of automatic and manual evaluation methods.
@@ 2 Related Work
Cultural knowledge, encompassing the customs, beliefs, traditions, and practices of a culture, is crucial yet challenging to encapsulate. While some researchers focus on manually curating cultural knowledge datasets, others evaluate LLMs’ performance on culturally related tasks. Yin et al. (2022) and Palta and Rudinger (2023) have developed benchmarks such as geo-diverse prompts and food-custom datasets (FORK) to probe cultural biases in commonsense reasoning systems. However, manual dataset construction is inefficient and hard to scale, prompting a shift towards automated methods. For instance, StereoKG Deshpande et al. (2022) offers a scalable knowledge graph that blends cultural knowledge with stereotypes, and Candle Nguyen et al. (2023) extracts cultural commonsense knowledge from the web, organizing it into clusters. Despite these advances, the variability in data representation—from sentences to triplets using OpenIE—poses challenges for consistency and noise control in knowledge probing. Keleg and Magdy (2023) aims to mitigate this by selecting culturally diverse factual triples from Wikidata, focusing mainly on explicit country information. In contrast, our work proposes an automated, efficient approach to constructing a cultural knowledge dataset in a uniform triplet format, addressing the limitations of existing methods and focusing on implicit cultural knowledge. Appendix A.1 shows a comparison of our dataset with existing ones.
@@
Deciphering the knowledge encoded by LLMs poses significant challenges due to their opaque nature , early benchmarks like LAMA Petroni et al. (2019) sought to quantify the factual knowledge in English LLMs, while ParaRel Elazar et al. (2021) highlighted their consistency issues. Subsequent efforts as mLAMA Kassner et al. (2021) and mParaRel Fierro and Søgaard (2022) expanded these benchmarks multilingually, though such methods often focus on single-word entities, limiting their depth of assessment. To address these shortcomings, newer studies Shin et al. (2020); Zhong et al. (2021); Meng et al. (2022) have evolved towards eliciting more comprehensive factual knowledge, including multi-word entities, with Jiang et al. (2020a) developing algorithms for multi-token predictions. LPAQA Jiang et al. (2020b) further refines this by optimizing prompt discovery for more accurate knowledge probing. Our work builds on this foundation, targeting multi-token probing within the food domain, characterized by complex expressions like Trigonella foenum-graecum.
@@ 3 FmLAMA Construction
iii. Properties included. The property “has part(s)” identifies food ingredients for a dish. Additional properties like “made from material” and “image” are collected to support future research (e.g., multimodal), though they are not utilized in this study. In each instance, the descriptive language for all attributes remains consistent.
Ultimately, our constructed dataset, FmLAMA, comprises 33,600 dish instances, detailed by name, origin, ingredients, and optionally, materials and images, encompassing 128 cultural country groups and 250 languages. The average number of ingredients in FmLAMA dataset is 2.04. Examples and statistics are provided in Appendix A.
@@ 4.1 Probing task
We adopt Word Prediction (WP) as the knowledge-probing task, following previous work Fierro and Søgaard (2022); Wu et al. (2023); Fierro et al. (2024b). Firstly, we manually design a prompt template $t$ focused on the core attribute “has part(s)”, illustrating a connection between a dish (subject) and its ingredient(s) (object). WP is usually implemented as a candidate retrieval problem. The candidate set consists of all objects in the focused, filtered sub-dataset of FmLAMA. [^2] As the size of the filtered sub-dataset increases, the candidate object set also expands, leading to greater difficulty in probing. Consequently, in this paper, results obtained by probing across different filtered sub-datasets cannot be used for horizontal comparison. The primary objective is to utilize LLMs to obtain the probability of each candidate $C$ and subsequently rank the predicted objects based on these probabilities.
@@
Using subject-object tuples ([X], [Y]) as queries, we probe LLMs by replacing the subject and masking the object. Considering that each candidate object is tokenized into $d$ subtokens $\left\{c_{1},\cdots,c_{d}\right\}$ by LLMs correspondingly, we apply [MASK] token of varying lengths to the objects within each query. $D$ queries are constructed for each dish subject $X$ based on the same template $t$, $D$ is the maximum number of object tokens, each query $Q_{d}^{X}$ is defined as:
$Q_{d}^{X}=t\left(X,\left[\mathrm{MASK}\right]*d\right),d\in\left[1,D\right]\vspace{-3pt}.$ (1)
@@
We use Mean Pooling [^3] Mean Pooling has been shown to be superior to other pooling methods for multi-token probing Wu et al. (2023). method to obtain the prediction probability of each candidate. Specifically, for candidate object $C=\left\{c_{1},\cdots,c_{d}\right\}$ of length $d$, its probability is obtained from the likelihoods associated with the [MASK] tokens in $Q_{d}^{X}$, and the predicted probability of $C$ is calculated as the average of the probabilities of composing its subtokens:
$P\left(Q_{d}^{X},C\right)=\frac{1}{d}\sum_{i=1}^{d}{p\left(Q_{d}^{X},\left[\mathrm{MASK}\right]_{i}=c_{i}\right)}\vspace{-3pt},$ (2)
where $p\left(\cdot\right)$ is obtained after the softmax operation.
4.2 Probing template
Considering LLMs produce varied predictions based on prompt framing Elazar et al. (2021); Wang et al. (2023), we craft five templates conveying identical meanings in each prompt language. For example, “[X] is a dish made with [Y]”, “[X] is a type of food that includes [Y]”.
Cultural information.
To explore the impact of introducing explicit cultural context on LLMs’ ability to access cultural knowledge, we enhance the basic templates by integrating country information. For instance, “In [C], [X] is a dish made with [Y]”, where [C] denotes the country of origin for the dish [X], [Y] indicates the ingredient object.
Multilingual prompts.
To explore how different prompt language settings affect LLMs’ cultural knowledge probing abilities, we craft prompts in six languages written by native speakers, including English (en), Chinese (zh), Arabic (ar), Korean (ko), Russian (ru), and Hebrew (he). These languages span 4 different language families – Indo-European (English, Russian), Semitic (Hebrew, Arabic), Altaic (Korean), and Sino-Tibetian (Chinese), and are spoken by more than 2.356 billion speakers. Furthermore, these languages represent cultural diversity, being spoken on different continents by groups with rich and distinct cultural backgrounds. These prompt illustrations and language details are depicted in Appendix B.
Code-switching.
[…]
@@ 4.3 Probing Metric
Given these constraints, we introduce two automatic metrics: the absolute-match metric, Mean Average Precision (mAP), and the fuzzy-match metric, Mean Word Similarity (mWS), along with a Manual Evaluation Score (MES) that incorporates both LLM-simulated and real human assessments for implementation. All mAP, mWS, and MES metrics are in the range $[0,1]$.
@@
mAP is widely used in information retrieval settings, assessing the relevance of predicted objects (in our case ingredients) only when they precisely match the reference ones. The precision at rank $k$ (P@k) for a given food instance $i$ is defined as follows:
@@
where $\mathrm{ing}_{i}$ is the reference ingredients set, and $\mathrm{topk}_{i}$ signifies the set of top-$k$ objects with the highest predicted probability of belonging to food item $i$ by LLMs. Then the average precision of food item $i$ is computed as follows:
$\mathrm{AP}_{i}=\frac{1}{\left|\mathrm{ing}_{i}\right|}\sum_{k=1}^{n}{\mathrm{P}@\mathrm{k}\times\mathrm{rel}@\mathrm{k}}\vspace{-3pt},$ (4)
where $n$ refers to the size of the candidate object set, and $\mathrm{rel}@\mathrm{k}$ is 1 if the object at rank $k$ is relevant to food item $i$, otherwise 0. Finally, we compute the mAP in the following way:
@@
where $G$ represents a dish group we are focusing on (i.e. a subset of FmLAMA).
@@
mWS is defined based on the semantic similarity between predicted and reference objects. First, we define the similarity score $S\left(i,g\right)$ for each ingredient $g$ within each dish $i$. Only the predicted objects in the top-$l$ rankings of the model prediction that are most similar to $g$ contribute to the evaluation score, where $l=\left|\mathrm{ing}_{i}\right|$. $S\left(i,g\right)$ is defined as follows:
@@
where $w_{g}$ and $w_{p}$ are the Fasttext Bojanowski et al. (2017); Joulin et al. (2016) vectors for the ingredient $g$ and the predicted object $p$, respectively, and $\cos\left(\cdot\right)$ is the cosine similarity function. Then we can compute the probing similarity $WS_{i}$ for each food instance $i$ and mWS for the targeted food group as follows: